# Novel Bounds From The Weak Gravity and Festina Lente Conjectures

### v1

@@ IV.1 The Higgs Instability
As we saw above, the magnetic form on the FLC suggests that there is a cutoff scale $\Lambda$ of the EFT associated with any gauged $U(1)$, and it would be interesting to see what this implies for the Higgs vacuum (in)stability. We first define $\lambda_{\text{eff}}\equiv 4V_{\text{eff}}/h^{4}$ as the effective Higgs quartic coupling. Then, the magnetic form of the FLC implies that
@@
If we interpret the scale at which the SM quartic supposedly runs into negative values as the scale of NP, such that $h=\Lambda$, and use the (field-dependent) mass of the electron as being the lightest particle in the SM QED, then we can express Eq. (20) as
@@
This bound is shown in Figure 2. If we use $\Lambda\sim 10^{9}$ GeV to be conservative, then we obtain $\lambda_{\text{eff}}\lesssim 4.4\times 10^{-4}$, whereas if we use $\Lambda\sim 10^{16}$ GeV as the upper limit on the scale on NP, then we obtain $\lambda_{\text{eff}}\lesssim 4.4\times 10^{-18}$. This limit is weaker than $\lambda_{\text{eff}}\lesssim O(10^{-22})$ obtained in [23].
IV.2 Inflation and milli-charged particles
Although the electric charge in the SM is quantized, with no experimental evidence so far to indicate any deviation of the charges of SM particles from quantization, it is nonetheless possible for the electric charge to deviate from the SM prediction. More specifically, the requirements of anomaly cancellation and the gauge invariance of the SM Yukawa interactions are alone insufficient to determine charge quantization, and they leave one degree of freedom that allows the electric charge in the SM to deviate from their values by a milli-charge parameter $\epsilon$. In addition, it is not hard to visualize the existence of particles with fractional charges depending on $\epsilon$ that are not quantized. Such particles are known as mCPs.
There many ways milli-charge can be introduced, including adding an $SU(2)\times SU(2)$ Dirac fermion with hypercharge $2\epsilon$ [16], introducing an additional $U(1)$ gauge group [17] which mixes with the SM $U(1)_{\text{em}}$ via kinetic mixing, utilizing the aforementioned remaining degree of freedom in the SM to allow neutrinos to have a milli-charge [18, 19], or through introducing non-locality in QED [20, 21]. Such proposal are interesting for other reasons, such as providing dark matter candidates like the dark photon, which arises in additional $U(1)$ scenarios where the gauge boson is stable or long-lived.
A simple and very straightforward application of the FLC in Eq. (3) is that it can be used to set constraints on the parameter space of mCPs. This was done in [22], where reasonably stringent bounds on the charge of mCPs, were obtained. Nonetheless, a major limitation on the applicability of the FLC for practical purposes lies in the smallness of the Hubble scale $H$ at the current epoch of the universe ($H_{0}$). However, this has not always been the case, and $H$ is expected to be much larger during the epoch of cosmological inflation, $H_{I}\gg H_{0}$. As the FLC is expected to apply in all epochs of the universe during which the vacuum was of dS type, particularly during inflation it reads [^5] More accurately and as pointed out in [30], the FLC should only apply in dS space where the vacuum energy is dominant, which implies that $\frac{\rho_{R}}{\rho_{V}}\geq c$, where $\rho_{R}$ is the radiation energy density and $c$ is a constant of $O(1)$.,
@@ IV.2 Inflation and milli-charged particles
As we will see below, applying the FLC during inflation enables us to obtain much stronger bounds than the naive application of the FLC at the current epoch. To show how this is done, recall that after the end of inflation the inflaton field decays causing reheating. Assuming that the decay is instantaneous, or at least fast enough so that almost all of the energy from the inflaton decay is converted into radiation, one finds
$3M_{P}^{2}H_{I}^{2}\simeq\frac{\pi^{2}}{30}g_{\text{reh}}T_{\text{reh}}^{4}\hskip 5.69054pt\implies T_{\text{reh}}=\Bigg(\frac{90M_{P}^{2}H_{I}^{2}}{\pi^{2}g_{\text{reh}}}\Bigg)^{\frac{1}{4}}\,$, (23)
where $g_{\text{reh}}$ is the number of effective degrees of freedom during reheating. For a successful Big Bang Nucleosynthesis (BBN), we need $T_{\text{reh}}\gtrsim T_{\text{BBN}}$, which from Eq. (23) implies a lower bound on the scale of inflation
@@
where we have used $T_{\text{BBN}}=1$ MeV, and $g_{\text{reh}}=10.75$. The bound is shown in Fig. 3, together with the other experimental bounds. In addition, we show the bound obtained by a naive application of the FLC as suggested by Eq. (3), which was obtained in Ref. [22]. As the plot shows, the bound from the inflationary FLC is orders of magnitude stronger than naive application of FLC. Note that this is a conservative bound on the scale of inflation assuming $T_{\text{reh}}\gtrsim T_{\text{BBN}}=1$ MeV. However, the actual reheating temperature may be much larger than $\sim 1$ MeV, which would help strengthen the bound even further.
@@ IV.3 The scale of inflation
This is the same result as we obtained earlier by demanding gravity to be the weakest force, as shown in Eq. 10. For the SM $U(1)_{\text{em}}$, this is trivially satisfied by the electron in the current epoch of the universe. However, it poses a serious problem for inflation once we set $H=H_{I}$, as it would imply an inflation scale being below the mass of the electron, which is not viable. This issue was raised in [12, 30] from the direct application of the FLC, however, unlike those results, Eq. (26) is independent of any $U(1)$ coupling and only depends on the mass of the lightest charged particle.
@@ V The FLC and Naturalness
where $\Lambda$ is the cutoff scale. Naturalness implies that $\delta m^{2}\lesssim m^{2}$. Thus, we can use $\delta m^{2}$ in the FLC instead of $m^{2}$. This implies a lower limit on $\Lambda$
@@
which implies a lower bound on the $U(1)$ charge $q\geq\sqrt{6}H/M_{P}$. This is identical to the result obtained in [13] from the combined application of the FL and the WGC. However, here we obtain the same result from the consideration of naturalness. If we use the current value of $H_{0}$, then we get a bound $q\gtrsim 10^{-60}$. However, we can obtain stronger bounds from inflation (see Eq. (24))
@@
however, this is much weaker than the current experimental bounds of $H_{I}<10^{13-14}$ GeV obtained from Planck+BICEP/Keck 2018 observation [32].
VI Conclusions
In this paper, we have studied several phenomenological implications of the FLC and the WGC. In particular, we have shown that the FLC and WGC can set stringent bounds on fifth force searches and could even exclude the entire parameter space for light particles with masses not much larger than the Hubble scale. We have also investigated the implications of the magnetic FLC for the Higgs potential instability, mCPs and the scale of inflation. In particular, we have found that the FLC sets stringent bounds on the Higgs quartic coupling. Moreover, when the FLC is applied during inflation, the requirement for the reheating temperature to be larger than the BBN temperature can set a stringent bound on the parameter space of mCPs. The latter turns out to be orders of magnitude stronger than what is suggested by a naive application of the FLC.
We have also shown that the magnetic FLC, as well as the assumption that gravity is the weakest force, can set an upper bound on the scale of inflation that depends only on the mass of the lightest charged particle, which ostensibly leads to tension between the FLC and inflation. However, this tension can be resolved if the lightest charged particle is allowed to become heavier in the UV. We have evaluated the implications of interplay between naturalness and the FLC/WGC, and found that this implies a lower limit on the charge of a $U(1)$ gauge group. We point out that the FLC can augment experimental searches in placing limits on the parameter space associated and any $U(1)$ gauge group, whether broken or unbroken.

### v2

@@ III.1 The Higgs Instability
As we saw above, the magnetic form of the FLC suggests that there is a cutoff scale $\Lambda$ on the EFT associated with any gauged $U(1)$, and it would be interesting to see what this implies for the Higgs vacuum (in)stability. We first define $\lambda_{\text{eff}}\equiv 4V_{\text{eff}}/h^{4}$ as the effective Higgs quartic coupling. Then, the magnetic form of the FLC implies that
@@
If we interpret the scale at which the SM quartic supposedly runs into negative values as the scale of NP [^6] This is a strong assumption, as it implicitly assumes no new physics up to that scale. Therefore, all the bounds derived here are based on this scenario., such that $h=\Lambda$, and use the (field-dependent) mass of the electron as being the lightest particle in the SM QED, then we can express Eq. (13) as
@@
This bound is shown in Figure 1. If we use $\Lambda\sim 10^{9}$ GeV to be conservative, then we obtain $\lambda_{\text{eff}}\lesssim 4.4\times 10^{-4}$, whereas if we use $\Lambda\sim 10^{16}$ GeV as the upper limit on the scale of NP, then we obtain $\lambda_{\text{eff}}\lesssim 4.4\times 10^{-18}$. This limit is weaker than $\lambda_{\text{eff}}\lesssim O(10^{-22})$ obtained in [24].
III.2 Inflation and milli-charged particles
Although the electric charge in the SM is quantized, with no experimental evidence so far to indicate any deviation of the charges of SM particles from quantization, it is nonetheless possible for the electric charges to deviate from the SM predictions. More specifically, the requirements of anomaly cancellation and the gauge invariance of the SM Yukawa interactions alone are insufficient to determine charge quantization, and they leave one degree of freedom that allows the electric charge in the SM to deviate from their values by a milli-charge parameter $\epsilon$. In addition, it is not hard to visualize the existence of particles with fractional charges depending on $\epsilon$ that are not quantized. Such particles are known as mCPs.
There are many ways a milli-charge can be introduced, including adding an $SU(2)\times SU(2)$ Dirac fermion with hypercharge $2\epsilon$ [25], introducing an additional $U(1)$ gauge group [26] that mixes with the SM $U(1)_{\text{em}}$ via kinetic mixing, utilizing the aforementioned remaining degree of freedom in the SM to allow neutrinos to have a milli-charge [27], or through introducing non-locality in QED [28, 29]. Such proposals are interesting for other reasons, such as providing dark matter candidates like the dark photon, which arises in additional $U(1)$ scenarios where the gauge boson is stable or long-lived.
A simple and very straightforward application of the FLC in Eq. (3) is that it can be used to set constraints on the parameter space of mCPs. This was done in [23], where reasonably stringent bounds on the charge of mCPs were obtained. Nonetheless, a major limitation on the applicability of the FLC for practical purposes, lies in the smallness of the Hubble scale $H$ at the current epoch of the universe ($H_{0}$). However, this has not always been the case, and $H$ is expected to have been much larger during the epoch of cosmological inflation, $H_{I}\gg H_{0}$. As the FLC is expected to apply during all epochs of the universe during which the vacuum was dS [^8] More accurately, and as pointed out in [30], the FLC should only apply in dS space where the vacuum energy is dominant, which implies that $\frac{\rho_{R}}{\rho_{V}}\leq c$, where $\rho_{R}$ is the radiation energy density and $c$ is a constant of $\mathcal{O}(1)$., particularly during inflation, it reads ,
@@ III.2 Inflation and milli-charged particles
As we will see below, applying the FLC during inflation enables us to obtain much stronger bounds than the naive application of the FLC at the current epoch. To show how this is done, recall that after the end of inflation, the inflaton decays and reheats the Universe. We define $T_{\rm reh}$ as the temperature at the onset of radiation domination. Assuming instantaneous (maximally efficient) reheating, i.e. that the energy density at the end of inflation is promptly converted into radiation, $\rho_{\rm end}\simeq\rho_{R}(T_{\rm reh})$, with $\rho_{\rm end}\simeq 3M_{P}^{2}H_{I}^{2}$ and $\rho_{R}(T)=\frac{\pi^{2}}{30}g_{\rm reh}T^{4}$, we obtain
$3M_{P}^{2}H_{I}^{2}\simeq\frac{\pi^{2}}{30}g_{\rm reh}T_{\rm reh}^{4}\quad\Rightarrow\quad T_{\rm reh}=\left(\frac{90\,M_{P}^{2}H_{I}^{2}}{\pi^{2}g_{\rm reh}}\right)^{1/4}\,.$ (16)
Here $g_{\rm reh}\equiv g_{*}(T_{\rm reh})$ is the effective number of relativistic degrees of freedom at reheating. For a successful Big Bang Nucleosynthesis (BBN), we need $T_{\text{reh}}\gtrsim T_{\text{BBN}}$, which from Eq. (16) implies a lower bound on the scale of inflation
@@
where we have used $T_{\text{BBN}}=1$ MeV, and $g_{\text{reh}}=10.75$. The bound is shown in Fig. 3, where we also superimpose representative laboratory, astrophysical and cosmological constraints on mCPs for comparison. As the figure illustrates, the inflationary FLC bound is most relevant in the ultra-light, very small-charge regime, where it provides a complementary constraint, while for larger masses the parameter space is typically dominated by the existing limits shown in the same plot. For reference, we also display the constraint from the naive present-epoch application of the FLC in Eq. (3), as discussed in Ref. [23]; the inflationary bound is parametrically stronger due to $H_{I}\gg H_{0}$. Finally, Eq. (18) is conservative since it uses the minimal reheating requirement $T_{\rm reh}\gtrsim T_{\rm BBN}\simeq 1~\mathrm{MeV}$; a higher reheating temperature would correspond to a larger $H_{I}$ and thus further strengthen the bound.
Before we conclude, we emphasize that Eq. (18) implicitly assumes that the mass and charge of the particle sourcing the Schwinger discharge (e.g. the mCP charge carrier) remain approximately constant from the end of inflation through reheating and down to BBN. More generally, if the charge carrier acquires a time-dependent (field-dependent) mass—for instance through a Higgs-like scalar $\phi$ whose background value evolves in time—then the relevant mass during inflation can differ from its late-time value. As a simple example one may have
$m(\phi)=y\,|\phi|\qquad\text{or}\qquad m^{2}(\phi)=m_{0}^{2}+y^{2}\phi^{2}\,$, (19)
so that, if $|\phi_{\rm inf}|>|\phi_{0}|$, one finds $m_{\rm inf}\equiv m(\phi_{\rm inf})>m(\phi_{0})$. Since the FLC bound scales as $q_{\max}\propto m^{2}$, a larger inflationary mass relaxes (weakens) the bound inferred from Eq.(18) when expressed in terms of late-time parameters. Once the field has settled at its minimum $\phi_{0}$ and the mass becomes time-independent, the bound should be applied using the asymptotic mass $m(\phi_{0})$, i.e. Eq.(18) applies again.
@@ III.3 The scale of inflation
For the SM $U(1)_{\text{em}}$, this is trivially satisfied by the electron in the current epoch of the universe. However, it poses a serious problem for inflation once we set $H=H_{I}$, as it would imply an inflation scale being below the mass of the electron, which is not viable. This issue was raised in [19, 30] from the direct application of the FLC, however, unlike those results, Eq. (20) is independent of any $U(1)$ coupling and only depends on the mass of the lightest charged particle.
@@ IV The FLC and Naturalness
where $\Lambda$ is the cutoff scale. Naturalness implies that $\delta m^{2}\lesssim m^{2}$. Thus, we can use $\delta m^{2}$ in the FLC instead of $m^{2}$. This implies a lower limit on $\Lambda$:
@@
which implies a lower bound on the $U(1)$ charge $q\geq\sqrt{6}H/M_{P}$. This is identical to the result obtained in [20] from the combined application of the FLC and the WGC. However, here we obtain the same result from the consideration of naturalness. If we use the current value of $H_{0}$, then we get a bound $q\gtrsim 10^{-60}$. However, we can obtain stronger bounds from inflation (see Eq. (17)) [^10] We should point out that all such bounds should be understood as lower limits only, as opposed to describing the phenomenology of actual particles.
@@
however, this is much weaker than the current experimental bounds of $H_{I}<10^{13-14}$ GeV obtained from Planck+BICEP/Keck 2018 observations [32].
V Conclusions
In this paper, we have studied several phenomenological implications of the FLC and the WGC. In particular, we have shown that the FLC and WGC can set stringent bounds on a fifth force. For fifth forces, the WGC/FLC interplay is most restrictive in the extreme-IR limit $(m\sim H)$; we present this limit as a parametric consistency check rather than a present-day exclusion. The phenomenologically relevant bounds we emphasize come from the inflationary application $((H\gg H_{0}))$ combined with reheating/BBN. We have also investigated the implications of the magnetic FLC for the Higgs potential instability, mCPs, and the scale of inflation. In particular, we have found that the FLC sets stringent bounds on the Higgs quartic coupling. Moreover, when the FLC is applied during inflation, the requirement for the reheating temperature to be larger than the BBN temperature can set a stringent bound on the parameter space of mCPs. The latter turns out to be orders of magnitude stronger than what is suggested by the naive application of the FLC.
We have also shown that the magnetic FLC, as well as the assumption that gravity is the weakest force, can set an upper bound on the scale of inflation that depends only on the mass of the lightest charged particle, which ostensibly leads to tension between the FLC and inflation. However, this tension can be resolved if the lightest charged particle is allowed to become heavier in the UV. We have evaluated the implications of the interplay between naturalness and the FLC/WGC, and found that this implies a lower limit on the charge of a $U(1)$ gauge group. We point out that the FLC can augment experimental searches in placing limits on the parameter space associated with any $U(1)$ gauge group, whether broken or unbroken.